# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ 4.2.2 Properties of the contributing MBHBs
We now investigate the properties of the MBHBs that contribute the most to the GWB across the PTA frequency band. In particular, we compute the contribution of binaries to the total characteristic strain as a function of their redshift and chirp mass. We consider a PTA experiment with observation span $T_{\rm obs}=T_{\rm EPTA}\simeq 10.33$ years. As mentioned above, the properties of the total characteristic GW strain vary significantly between the low and high frequency end of the PTA sensitivity band, so we take two representative frequencies: $1/T_{\rm EPTA}\simeq 3.1\,\rm nHz$ and $9/T_{\rm EPTA}\simeq 27.6\,\rm nHz$. In Figure 3, we show the differential contribution of the MBHB population divided into redshift slices. The solid line represents the median values across universe realizations for each redshift slice, while the shaded region indicates the 68% confidence interval, defined here by the 16th and 84th percentiles of the distributions. At low frequencies, for circular and eccentric populations, the peak of the distribution is found around $z\simeq 0.3-0.4$. At lower redshifts, we observe a larger variance in the contribution compared to higher redshifts. Since the variance (width of the confidence interval) is primarily determined by the average number of sources contributing to the GW strain in each bin, this reflects the smaller number of MBHBs expected at low redshift. At high frequency, the peak of distribution for both populations shifts towards a higher redshift, around $z\simeq 0.6-0.7$. The heavy nearby systems, which dominate the GW signal at low frequencies, spend very little time at high frequencies; therefore, the signal accumulates over a larger volume. This also explains the very large variance compared to the low-frequency distribution: we have a lower number of binaries that contribute at high frequencies. Again, at high frequencies, the eccentric population exhibits smaller variance due to the increased number of contributing binaries. The typical redshift ranges of the MBHBs contributing to 90% of the total characteristic strain signal are given by the vertical lines in Figure 3. These correspond to the median redshift values (across universe realizations) at which the cumulative contribution — integrated from low and high redshifts, respectively — reaches 5% of the total strain signal. For the low frequency of 3.1 nHz considered here, where the GW signal is expected to be seen as a stochastic background, we find that most of the contribution comes from the range $0.05<z<1.2$ for the circular population and $0.08<z<1.3$ for the eccentric population.
@@
Similar distributions as a function of MBHB chirp mass are shown in Figure 4. These distributions confirm what we have mentioned above. The GW signal at low frequencies is mainly formed by heavy systems with chirp mass in the range $10^{8.5}-10^{9.5}\,{\rm M_{\odot}}$, while most of the contribution at high frequencies comes from binaries in the range $10^{8}-10^{9}\,{\rm M_{\odot}}$. At low frequency, we find that the contribution to the total strain signal extends to higher masses for the eccentric population relative to the circular population. The main reason is the contribution of the massive and eccentric binaries with orbital period around and even below the $0.5/T_{\rm obs}$ frequency, emitting at higher orbital harmonics. The typical lower bound on the chirp mass of MBHBs contributing 90% of the total characteristic strain signal is indicated by the vertical lines in Figure 4. These represent the median chirp mass values (across universe realizations) at which the cumulative contribution — integrated from the lowest masses — reaches 10% of the total strain signal, for each population. We find $\log_{10}(\mathcal{M}_{\rm c}/\rm M_{\odot})\gtrsim 8.35$ for the circular population and $\log_{10}(\mathcal{M}_{\rm c}/{\rm M_{\odot}})\gtrsim 8.63$ for the eccentric population.
@@ 4.2.3 Number of contributing MBHBs
where the summation in the numerator is performed over the merger population in decreasing order of $h_{\rm c}^{2}$. We use Equation 30 to compute each binary’s contribution. In this analysis, we adopt $\alpha_{\rm tot}=0.75$, which corresponds to a signal-to-noise ratio (SNR) of 3 for the brightest sources relative to the rest of the MBHB population, defined as $\sum_{l=1}^{N}h_{{\rm c},l}^{2}(f)/\sum_{l\geq N+1}h_{{\rm c},l}^{2}(f)=3$.
However, this metric does not quantify the relative contribution of the ”brightest sources”. For example, assume that 74% of the GWB power is produced by a few binaries and the remaining 1% needed to reach 75% comes from 999 MBHBs, in this case we would obtain $N_{75}=1000$ even though the GWs from the population is dominated by one system. To capture this potential problem, we suggest another measure, $N_{\rm eff}$, inspired by the effective sample size of a Markov Chain, derived in Elvira et al. (2018), which can be written as
@@
The (effective) number of binaries that contribute the most to the total GW strain across the PTA frequency band is presented in Figure 5. The results for $N_{75}$ are given as error boxes and median $N_{\rm eff}$ values are presented by circles for the circular population and crosses for the eccentric population, for our 2 0002\;000 universe realizations. Below 10 nHz, for both populations, the typical number of binary signals that must be added to reach 75% of the background signal is greater than 300, indicating that the stochastic approximation seems justifiable at those frequencies. However, around and above $f_{\mathrm{yr}}$, for 25% of the universe realizations, $N_{75}$ is lower than 10, which is not sufficiently high to apply the central limit theorem. The eccentric background has a factor $\approx 10$ greater number of contributors ($N_{75}$) around $f_{\mathrm{yr}}$ due to the distributed spectrum of GW emission and therefore maintains stochasticity to higher frequencies.
@@
At higher frequencies, the contributions are more balanced and point to a number of contributing binaries from few to tens binaries. This implies that at higher frequency, the background is less likely to be well approximated by a Gaussian noise and therefore is less similar to the conventional definition of a background. Taking the frequency range as a whole, one might conclude that detecting individual sources is unlikely due to the large number of contributing binaries, especially at low frequencies. However, even if the fractional contribution of an individual source is small, the cumulative effect of observing multiple pulsars within a given PTA can enhance its detectability. Furthermore, the detectability of an individual binary by a PTA must consider both pulsar noise and the detector response to GWs. This is examined in the next section, where we investigate how to pass from the properties of the strain signal to the PTA observables.
@@ 5 Implications for Pulsar Timing Array observations
To accurately compare the theoretical signal computed from the Horizon-AGN population with the one currently measured by PTA experiments, we must study the properties of the GWB-induced timing residuals. Timing residuals are the time difference between the measured time of arrivals (ToAs) of radio pulses from Galactic millisecond pulsars and the ToAs predicted by a theoretical model, referred to as the timing model (Edwards et al., 2006). The residuals also carry an imprint from GWs; in particular, a GWB would appear as a common correlated signal with a spatial correlation pattern first derived in Hellings & Downs (1983).
@@ 5.1 Methodology
In this section, we present two methods for estimating the spectral properties of the timing residuals induced by the MBHB population. We emphasize the importance of considering what affects the spectral inference of the detector (here the PTA), to avoid potential biases when interpreting its results. We also outline our methodology for identifying potentially resolvable individual binaries within the population.
@@ 5.1.1 The Gaussian ensemble approach
This approach considers the GWB signal generated by a MBHBs population as a Gaussian ensemble (see e.g. Rosado et al., 2015; Romano & Cornish, 2017; Hazboun et al., 2019; Agazie et al., 2023c; EPTA Collaboration et al., 2024a). In this case, the GW strain is considered as a random variable, fully characterized by its one-sided power spectral density (PSD), $S_{\rm h}(f)=h_{\rm c}^{2}(f)/f$. If one assumes that the GWB signal is stationary, isotropic and un-polarized, the average timing residuals response is given, in terms of PSD, by (Bertotti et al., 1983; Jenet et al., 2006)
@@
The timing residual PSDs from the different universe realizations obtained using this method have two limitations. First, they do not account for the fact that the signal is produced by a discrete number of sources, each with its own response to the PTA. Second, the derivation of $h_{\rm c}^{2}$ that leads to the timing residuals PSD $S_{\rm r}^{(\mathrm{GWB})}$, assumes that each MBHBs contributes a Dirac delta function at its orbital frequency harmonics as in Equation 30. We recall below that this is only the case in the limit of an infinitely long dataset [^6] It also assumes that the GW sources exhibit no frequency evolution.. For realistic PTA data, the effect of the finite observation duration on spectral estimation needs to be considered a priori. We use a second method for evaluating $S_{\rm r}^{(\mathrm{GWB})}$ that accounts for both of these issues in the following subsection.
@@ 5.1.2 The population approach
The principle of this approach is to compute the GWB-induced timing residuals summing the contributions of each individual MBHB (see e.g. Bécsy et al., 2022). This requires knowledge of the waveform for any set of orbital parameters. For this approach, we consider only circular binaries.
The exact residuals induced for a given pulsar by an inspiralling MBHB depend on their relative position, the binary inclination angle $\iota$ to the line of sight and the polarization angle $\psi$. The time series of timing residuals is composed of two terms, corresponding to the impact of the GW strain at the time of emission of the radio pulses (referred to as the Pulsar term) and at the time of their reception (referred to as the Earth term) (Detweiler, 1979). The correlation between GWB timing residuals among pulsars is entirely due to the Earth term. Thus, in the following, we only consider the GW strain signal of the Earth term when computing the GWB timing residuals power spectrum. We assume GWB to be isotropic, reflecting the distribution of MBHBs across the sky. As a result, we expect the GWB to imprint, on average, the same signal in every pulsar (Hellings & Downs, 1983); therefore, we compute the GWB spectrum for only one pulsar and assume that it is the same for the other pulsars in the array.
@@
Contrary to the previous case where we considered Dirac delta functions as a contribution of each MBHB, the finite width of the sinc-function leads to spectral leakage in the non-white PSD. As a result, every MBHB can potentially contribute to several frequencies, which makes frequency bins correlated. It is common practice in signal processing to taper the time series to minimize the effect of spectral leakage (Percival & Walden, 1993). However, in current PTA data analysis, no such pre-processing is applied to the timing residuals time series, so it might, a priori, affects the inference of the PSD. This suggests that even MBHBs emitting GWs at frequencies much lower than the PTA frequency ($1/T_{\rm obs}$) could contribute to the GWB power inferred in the PTA band due to spectral leakage.
The spectral leakage could (at least partially) be mitigated by the timing model. A part of the timing model fits and removes linear and quadratic trends in the PTA residuals that could be caused by the spin-down in the millisecond pulsars. Obviously, it also removes some very low frequency GWs covariant with such a trend (Cordes & Shannon, 2010; van Haasteren & Levin, 2013). As a result, it can significantly reduce the dynamical range in the expected PSD and decrease the effect of spectral leakage. An effect already discussed in Taylor et al. (2013); Lentati et al. (2013). However, the timing model cannot be a replacement for a well designed taper that controls the leakage. Nonetheless, the timing model acts as a high-pass filter which removes very low-frequency GWB ($\lesssim 1/T_{\rm obs}$) from the pulsar timing residuals (Hazboun et al., 2019). As a result, we consider only MBHBs that emit GWs with $f_{\mathrm{GW}}>1/(2T_{\rm obs})$ in order to reduce the numerical cost of the GWB computation without affecting its inferred properties in the PTA frequency band.
We produce realistic GWB timing residuals by summing the contributions from all MBHBs of a given universe realization, uniformly drawing the sky location, polarization angle, and initial phase for each binary. However, since the number of MBHBs contributing to the PTA band typically exceeds millions, we employ a numerical simplification allowing us to do this efficiently; the procedure is described in Appendix B. Once the GWB timing residuals waveform $\tilde{r}_{\rm GWB}$ is obtained, we can return to the time domain and infer the GWB PSD using the data analysis procedure commonly employed in the PTA community. In the following, we briefly discuss the PSD inference procedure used for PTA data.
@@ 5.1.3 The GWB spectral inference in PTA
This defines a prior for the noise Fourier coefficients, that depends on some parameters $\theta_{X}$, which characterize the PSD of the process. It is then usual procedure to marginalize over these coefficients so that the PTA likelihood depends only on the PSD parameters $\theta_{X}$ (Lentati et al., 2013; van Haasteren & Vallisneri, 2014). The observation time $T_{\rm obs}$, in Equation 40, is defined depending on whether the noise component, $X$, is intrinsic to a specific pulsar or not. If it is intrinsic, $T_{\rm obs}$ is the observation duration of that pulsar. If $X$ is a common process, such as a GWB, $T_{\rm obs}$ is the total observation duration of the PTA.
@@
which implies a power-law form $S_{\rm r}^{(\mathrm{GWB})}\propto f^{-\gamma}$ for the timing residuals PSD with $\gamma=3-2\alpha$, using Equation 35. The spectral index $\gamma$ and an amplitude at fiducial frequency $f_{\mathrm{yr}}=1/\rm{yr}$ are inferred from the data using a Bayesian framework. This requires a choice of prior for the GWB parameters $\theta_{\rm GWB}=\left(A_{\rm yr},\gamma\right)$. We use log-uniform prior for the GWB amplitude, $\log_{10}A_{\mathrm{yr}}\in[-18,-10]$ and uniform prior for its spectral index, $\gamma\in[0,7]$.
@@
In the following, we apply this PTA Bayesian spectral estimation method to the GWB-induced timing residuals of a single pulsar, obtained using the population approach described above. Since we are only interested in the accuracy of the spectral inference and its potential limitations (e.g., regarding spectral leakage), we simplify simulated PTA data and include/consider only white noise residuals in addition to the GWB timing residuals associated with each universe realization [^7] To obtain the same SNR for the GWB despite using only one pulsar, we scale its white noise residuals RMS as $\sigma_{\rm WN}=\sigma_{\rm PTA}/\sqrt{N_{\rm psr}}$, where $\sigma_{\rm PTA}$ denotes the characteristic ToA measurement error of the considered PTA made of $N_{\rm psr}$ pulsars..

### v2

@@ 4.2.2 Properties of the contributing MBHBs
We now investigate the properties of the MBHBs that contribute the most to the GWB across the PTA frequency band. In particular, we compute the contribution of binaries to the total characteristic strain as a function of their redshift and chirp mass. We consider a PTA experiment with observation span $T_{\rm obs}=T_{\rm EPTA}\simeq 10.33$ years. As mentioned above, the properties of the total characteristic GW strain vary significantly between the low and high frequency end of the PTA sensitivity band, so we take two representative frequencies: $1/T_{\rm EPTA}\simeq 3.1\,\rm nHz$ and $9/T_{\rm EPTA}\simeq 27.6\,\rm nHz$. In Fig. 3, we show the differential contribution of the MBHB population divided into redshift slices. The solid line represents the median values across Universe realizations for each redshift slice, while the shaded region indicates the 68% confidence interval, defined here by the 16th and 84th percentiles of the distributions. At low frequencies, for circular and eccentric populations, the peak of the distribution is found around $z\simeq 0.3-0.4$. At lower redshifts, we observe a larger variance in the contribution compared to higher redshifts. Since the variance (width of the confidence interval) is primarily determined by the average number of sources contributing to the GW strain in each bin, this reflects the smaller number of MBHBs expected at low redshift. At high frequency, the peak of distribution for both populations shifts toward a higher redshift, around $z\simeq 0.6-0.7$. The heavy nearby systems, which dominate the GW signal at low frequencies, spend very little time at high frequencies; therefore, the signal accumulates over a larger volume. This also explains the very large variance compared to the low-frequency distribution: we have a lower number of binaries that contribute at high frequencies. Again, at high frequencies, the eccentric population exhibits smaller variance due to the increased number of contributing binaries. The typical redshift ranges of the MBHBs contributing to 90% of the total characteristic strain signal are given by the vertical lines in Fig. 3. These correspond to the median redshift values (across Universe realizations) at which the cumulative contribution — integrated from low and high redshifts, respectively — reaches 5% of the total strain signal. For the low frequency of 3.1 nHz considered here, where the GW signal is expected to be seen as a stochastic background, we find that most of the contribution comes from the range $0.05<z<1.2$ for the circular population and $0.08<z<1.3$ for the eccentric population.
@@
Similar distributions as a function of MBHB chirp mass are shown in Fig. 4. These distributions confirm the findings noted above. The GW signal at low frequencies is mainly formed by heavy systems with chirp mass in the range $10^{8.5}-10^{9.5}\,{\rm M_{\odot}}$, while most of the contribution at high frequencies comes from binaries in the range $10^{8}-10^{9}\,{\rm M_{\odot}}$. At low frequency, we find that the contribution to the total strain signal extends to higher masses for the eccentric population relative to the circular population. The main reason is the contribution of the massive and eccentric binaries with orbital periods around and even below the $0.5/T_{\rm obs}$ frequency, emitting at higher orbital harmonics. The typical lower bound on the chirp mass of MBHBs contributing 90% of the total characteristic strain signal is indicated by the vertical lines in Fig. 4. These represent the median chirp mass values (across Universe realizations) at which the cumulative contribution (integrated from the lowest masses) reaches 10% of the total strain signal, for each population. We find $\log_{10}(\mathcal{M}_{\rm c}/\rm M_{\odot})\gtrsim 8.35$ for the circular population and $\log_{10}(\mathcal{M}_{\rm c}/{\rm M_{\odot}})\gtrsim 8.63$ for the eccentric population.
@@ 4.2.3 Number of contributing MBHBs
where the summation in the numerator is performed over the merger population in decreasing order of $h_{\rm c}^{2}$. We use Eq. 30 to compute each binary’s contribution. In this analysis, we adopt $\alpha_{\rm tot}=0.75$, which corresponds to a S/N of 3 for the brightest sources relative to the rest of the MBHB population, defined as $\sum_{l=1}^{N}h_{{\rm c},l}^{2}(f)/\sum_{l\geq N+1}h_{{\rm c},l}^{2}(f)=3$.
However, this metric does not quantify the relative contribution of the ”brightest sources”. For example, assume that 74% of the GWB power is produced by a few binaries and the remaining 1% needed to reach 75% comes from 999 MBHBs, in this case, we would obtain $N_{75}=1000$ even though the GWs from the population are dominated by one system. To capture this potential problem, we suggest another measure, $N_{\rm eff}$, inspired by the effective sample size of a Markov Chain, derived in Elvira et al. (2022), which can be written as
@@
The (effective) number of binaries that contribute the most to the total GW strain across the PTA frequency band is presented in Fig. 5. The results for $N_{75}$ are given as error boxes and median $N_{\rm eff}$ values are presented by circles for the circular population and crosses for the eccentric population, for our 2 0002\;000 Universe realizations. Below 10 nHz, for both populations, the typical number of binary signals that must be added to reach 75% of the background signal is greater than 300, indicating that the stochastic approximation seems justifiable at those frequencies. However, around and above $f_{\mathrm{yr}}$, for 25% of the Universe realizations, $N_{75}$ is lower than 10, which is not sufficiently high to apply the central limit theorem. The eccentric background has a factor $\approx 10$ greater number of contributors ($N_{75}$) around $f_{\mathrm{yr}}$ due to the distributed spectrum of GW emission and therefore maintains stochasticity to higher frequencies.
@@
At higher frequencies, the contributions are more balanced and point to a number of contributing binaries from a few up to several dozen binaries. This implies that at higher frequency, the background is less likely to be well approximated by a Gaussian noise and therefore is less similar to the conventional definition of a background. Taking the frequency range as a whole, one might conclude that detecting individual sources is unlikely due to the large number of contributing binaries, especially at low frequencies. However, even if the fractional contribution of an individual source is small, the cumulative effect of observing multiple pulsars within a given PTA can enhance its detectability. Furthermore, the detectability of an individual binary by a PTA must consider both pulsar noise and the detector response to GWs. This is examined in the next section, where we investigate how to pass from the properties of the strain signal to the PTA observables.
@@ 5 Implications for pulsar timing array observations
To accurately compare the theoretical signal computed from the Horizon-AGN population with the one currently measured by PTA experiments, we must study the properties of the GWB-induced timing residuals. Timing residuals are the time difference between the measured times of arrival (ToAs) of radio pulses from Galactic millisecond pulsars and the ToAs predicted by a theoretical model, referred to as the timing model (Edwards et al., 2006). The residuals also carry an imprint from GWs; in particular, a GWB would appear as a common correlated signal with a spatial correlation pattern first derived in Hellings and Downs (1983).
@@ 5.1 Methodology
In this section we present two methods for estimating the spectral properties of the timing residuals induced by the MBHB population. We emphasize the importance of considering what affects the spectral inference of the detector (here the PTA), to avoid potential biases when interpreting its results. We also outline our methodology for identifying potentially resolvable individual binaries within the population.
@@ 5.1.1 The Gaussian ensemble approach
This approach considers the GWB signal generated by an MBHBs population as a Gaussian ensemble (see e.g., Rosado et al., 2015; Romano and Cornish, 2017; Hazboun et al., 2019; Agazie et al., 2023c; EPTA Collaboration et al., 2024a). In this case, the GW strain is considered as a random variable, fully characterized by its one-sided power spectral density (PSD), $S_{\rm h}(f)=h_{\rm c}^{2}(f)/f$. If one assumes that the GWB signal is stationary, isotropic and un-polarized, the average timing residuals response is given, in terms of PSD, by (Bertotti et al., 1983; Jenet et al., 2006)
@@
The timing residual PSDs from the different Universe realizations obtained using this method have two limitations. First, they do not account for the fact that the signal is produced by a discrete number of sources, each with its own response to the PTA. Second, the derivation of $h_{\rm c}^{2}$ that leads to the timing residuals PSD $S_{\rm r}^{(\mathrm{GWB})}$, assumes that each MBHB contributes a Dirac delta function at its orbital frequency harmonics as in Eq. 30. We recall below that this is only the case in the limit of an infinitely long dataset [^6] It also assumes that the GW sources exhibit no frequency evolution.. For realistic PTA data, the effect of the finite observation duration on spectral estimation needs to be considered a priori. We use a second method for evaluating $S_{\rm r}^{(\mathrm{GWB})}$ that accounts for both of these issues in the following subsection.
@@ 5.1.2 The population approach
The principle of this approach is to compute the GWB-induced timing residuals by summing the contributions of each individual MBHB (see e.g. Bécsy et al., 2022). This requires knowledge of the waveform for any set of orbital parameters. For this approach, we consider only circular binaries.
The exact residuals induced for a given pulsar by an inspiralling MBHB depend on their relative position, the binary inclination angle $\iota$ to the line of sight and the polarization angle $\psi$. The time series of timing residuals is composed of two terms, corresponding to the impact of the GW strain at the time of emission of the radio pulses (referred to as the Pulsar term) and at the time of their reception (referred to as the Earth term; Detweiler 1979). The correlation between GWB timing residuals among pulsars is entirely due to the Earth term. Thus, in the following, we only consider the GW strain signal of the Earth term when computing the GWB timing residuals power spectrum. We assume GWB to be isotropic, reflecting the distribution of MBHBs across the sky. As a result, we expect the GWB to imprint, on average, the same signal in every pulsar (Hellings and Downs, 1983); therefore, we compute the GWB spectrum for only one pulsar and assume that it is the same for the other pulsars in the array.
@@
Contrary to the previous case where we considered Dirac delta functions as a contribution of each MBHB, the finite width of the sinc function leads to spectral leakage in the non-white PSD. As a result, every MBHB can potentially contribute to several frequencies, which makes frequency bins correlated. It is common practice in signal processing to taper the time series to minimize the effect of spectral leakage (Percival and Walden, 1993). However, in current PTA data analysis, no such pre-processing is applied to the timing residuals time series, so it might, a priori, affects the inference of the PSD. This suggests that even MBHBs emitting GWs at frequencies much lower than the PTA frequency ($1/T_{\rm obs}$) could contribute to the GWB power inferred in the PTA band due to spectral leakage.
The spectral leakage could (at least partially) be mitigated by the timing model. A part of the timing model fits and removes linear and quadratic trends in the PTA residuals that could be caused by the spin-down in the millisecond pulsars. Obviously, it also removes some very low-frequency GWs covariant with such a trend (Cordes and Shannon, 2010; van Haasteren and Levin, 2013). As a result, it can significantly reduce the dynamical range in the expected PSD and decrease the effect of spectral leakage. An effect already discussed in Taylor et al. (2013); Lentati et al. (2013). However, the timing model cannot be a replacement for a well-designed taper that controls the leakage. Nonetheless, the timing model acts as a high-pass filter that removes very low-frequency GWB ($\lesssim 1/T_{\rm obs}$) from the pulsar timing residuals (Hazboun et al., 2019). As a result, we consider only MBHBs that emit GWs with $f_{\mathrm{GW}}>1/(2T_{\rm obs})$ in order to reduce the numerical cost of the GWB computation without affecting its inferred properties in the PTA frequency band.
We produce realistic GWB timing residuals by summing the contributions from all MBHBs of a given Universe realization, uniformly drawing the sky location, polarization angle, and initial phase for each binary. However, since the number of MBHBs contributing to the PTA band typically exceeds millions, we employed a numerical simplification, which allowed us to carry out this procedure more efficiently; the procedure is described in Appendix B. Once the GWB timing residuals waveform $\tilde{r}_{\rm GWB}$ is obtained, we can return to the time domain and infer the GWB PSD using the data analysis procedure commonly employed in the PTA community. In the following, we briefly discuss the PSD inference procedure used for PTA data.
@@ 5.1.3 The GWB spectral inference in PTA
This defines a prior for the noise Fourier coefficients, that depends on some parameters $\theta_{X}$, which characterize the PSD of the process. It is then usual procedure to marginalize over these coefficients so that the PTA likelihood depends only on the PSD parameters $\theta_{X}$ (Lentati et al., 2013; van Haasteren and Vallisneri, 2014). The observation time $T_{\rm obs}$, in Eq. 40, is defined depending on whether the noise component, $X$, is intrinsic to a specific pulsar or not. If it is intrinsic, $T_{\rm obs}$ is the observation duration of that pulsar. If $X$ is a common process, such as a GWB, $T_{\rm obs}$ is the total observation duration of the PTA.
@@
which implies a power-law form $S_{\rm r}^{(\mathrm{GWB})}\propto f^{-\gamma}$ for the timing residuals PSD with $\gamma=3-2\alpha$, using Eq. 35. The spectral index $\gamma$ and an amplitude at fiducial frequency $f_{\mathrm{yr}}=1/\rm{yr}$ are inferred from the data using a Bayesian framework. This requires a choice of prior for the GWB parameters $\theta_{\rm GWB}=\left(A_{\rm yr},\gamma\right)$. We use a log-uniform prior for the GWB amplitude, $\log_{10}A_{\mathrm{yr}}\in[-18,-10]$ and a uniform prior for its spectral index, $\gamma\in[0,7]$.
@@
In the following, we apply this PTA Bayesian spectral estimation method to the GWB-induced timing residuals of a single pulsar, obtained using the population approach described above. Since we are only interested in the accuracy of the spectral inference and its potential limitations (e.g., regarding spectral leakage), we simplify simulated PTA data and include/consider only white noise residuals in addition to the GWB timing residuals associated with each Universe realization [^7] To obtain the same S/N for the GWB despite using only one pulsar, we scale its white noise residuals RMS as $\sigma_{\rm WN}=\sigma_{\rm PTA}/\sqrt{N_{\rm psr}}$, where $\sigma_{\rm PTA}$ denotes the characteristic ToA measurement error of the considered PTA made of $N_{\rm psr}$ pulsars..